OpenStudy (anonymous):

Using trapezoid VWYZ, solve for x. 29.3 14.3 6 15

OpenStudy (anonymous):

see the two angles are equal, as the sides ZV=YW so 4x+3=x+15 3x=12 x=4

OpenStudy (anonymous):

I dont have those numbers in my problem so would I put mine instead

OpenStudy (anonymous):

yes.